• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name S. Gilad

    Log entry time 16:04:42 on December 22,1999

    Entry number 31828

    keyword=morning shift 12/22

    Running smoothly . Short period of beam position instability during run 1550. Was fixewd by MCC.
    11:05am - DAQ crash. Have to reset ROC14.
    Took about 1 hour of coincidence, kinematics 16.
    Took about 40 minutes of dummy data.
    Problems in replay of coinc data - cannot figure out how many events we have. Decide to take another 1 hour of 3He data to beon the safe side.
    Replay problem was due to uncorrected raster. Have enough 3He events.
    12:05 - Replace target by C target.
    e-p check their DAQ system. Still broken.
    12:30-1:00 - C pointing run. End of Kinematics 16.
    Change electron arm to 14.22 degrees (14deg +69.2). electron momentum to 37711 ->3614.77 MeV (recycled Q2 and Q3).
    Take cross-calibration run with 3He.
    Take C pointing run at cross-calibration setting. Use run 1566 for good data.
    15:35pm. - move hadron arm to 67.83deg. set field to 680 -> 683.74 MeV. Loaded correct trigger file (a-la ROM).